OpenStudy (anonymous):

Determine the value of y for the system of equations. X-3y=3 3x+y=6

OpenStudy (anonymous):

multiply the first equation by -3 and add this gives an equation in y only which is easily solved

OpenStudy (anonymous):

need further help?

OpenStudy (anonymous):

Yha

OpenStudy (anonymous):

ok after multiplying first equation by -3 we get -3x + 9y = -9 every term in equation is multiplied by -3 3x + y = 6 add: 0 + 10y = -3 y = -3/10 or -0.3

OpenStudy (anonymous):

Thank u